OpenStudy (anonymous):

a) A flexible garden hose is coiled for storage. Each subsequent loop is 6 inches longer than the preceding loop, and the innermost loop is 34 inches long. If there are 6 loops, how long is the hose? b) Melinda makes monthly car payments of $285 each month. How much will she have paid after 2 years? 5 years? - total paid at 2 years? - total paid at 5 years? c) A display of soup cans is arranged with 1 can on top and each row having an additional can. How many cans are in a display of 20 rows?

OpenStudy (sleepyhead314):

Hello @alexneedshelpbad and Welcome to OpenStudy! :) a) first = 34 inches second = 34 + 6 = 40 inches third = 40 + 6 = 46 inches fourth = 46 + 6 = 52 inches fifth = 52 + 6 = 58 inches sixth = 58 + 6 = 64 inches so 34 + 40 + 46 + 52 + 58 + 64 = (total length)

OpenStudy (sleepyhead314):

@alexneedshelpbad You've closed this question post, do you still need help with the other parts to this question? :)

OpenStudy (anonymous):

no, i'm good! thank you :)